# Attention-based Reinforcement Learning for Combinatorial Optimization: Application to Job Shop Scheduling Problem

### 29 Jan 2024 v1

@@ Abstract
Job shop scheduling problems are one of the most important and challenging combinatorial optimization problems that have been tackled mainly by exact or approximate solution approaches. However, finding an exact solution can be infeasible for real-world problems, and even with an approximate solution approach, it can require a prohibitive amount of time to find a near-optimal solution, and the found solutions are not applicable to new problems in general. To address these challenges, we propose an attention-based reinforcement learning method for the class of job shop scheduling problems by integrating a policy gradient reinforcement learning with a modified transformer architecture. An important result is that our trained learners in the proposed method can be reused to solve large-scale problems not used in training and demonstrate that our approach outperforms the results in recent studies and widely adopted heuristic rules.
@@ 1 Introduction
The job shop scheduling problem (JSSP) is a well-known and one of the hardest classes of combinatorial optimization problems in operations research, computer science, and industrial engineering. It has a long history as a research topic and is still being actively studied because of its applicability in numerous industrial areas, such as manufacturing and production domains. (Gao et al.,, 2020).
@@
JSSP has been tackled in mainly two ways, by exact solution approaches and approximate approaches (Zhang et al.,, 2019; Kwon et al.,, 2021; Xie et al.,, 2019). Exact solution approaches with mathematical programming approaches have been widely adopted (Ku and Beck,, 2016; Pei et al.,, 2020; Wang et al.,, 2018; Zhang and Wang,, 2018), however, it may take a prohibitive amount of time to find an optimal solution, or the problems can be infeasible to find an optimal solution. This can be problematic for real-world JSSP, where a large number of jobs or machines would be associated. This has led researchers and practitioners to turn their attention to approximate methods, which aim to achieve near-optimal solutions instead. Various heuristic approaches have been applied to solve JSSP.(Gonçalves et al.,, 2005; Park et al.,, 2003; Peng et al.,, 2015; Kawaguchi and Fukuyama,, 2016; Chakraborty and Bhowmik,, 2015; Cruz-Chávez et al.,, 2017; Nagata and Ono,, 2018; Shi et al.,, 2020)
@@
Although the above mentioned approaches have gained success in addressing JSSP and are recognized as effective means to solve this particular problem type, solutions by those methods are subject to the constraints that are assumed when solving the problems. For example, solutions by the above methods are for the particular problems where the methods are applied, and thus, the solutions are not applicable to other problems in general. New problems are solved anew even with the same approach. Considering the classical solution approaches’ shortcomings it is essential to have a framework that can effectively solve large-scale problems without the need for model adjustments or human intervention.
@@
To address those challenges, we propose a learning method that can be reused to provide solutions to unseen problems. In particular, we employ attention mechanisms (Bahdanau et al.,, 2014) and integrate them with reinforcement learning (Sutton and Barto,, 2018) and the transformer architecture (Vaswani et al.,, 2017). We show that with our method, trained learners can be reused to solve large-scale problems not used in training and demonstrate that our proposed approach outperforms the results in recent studies and widely adopted heuristic rules.
@@ 2 Related Work
Vinyals et al. (Vinyals et al.,, 2015) proposed the Pointer Network, a sequence-to-sequence model to generate a permutation of the input sequence instead a generated sequence from a fixed-size dictionary. This allows the model to select and order elements from the input sequence. In the original work of Vinyals et al. (Vinyals et al.,, 2015), the traveling salesman problem (TSP) was studied. However, the model is trained with supervised learning and optimal solutions obtained from alternative methods, a notable limitation is which may not be practical in all cases. To address this issue Bello et al. (Bello et al.,, 2017) introduced a Reinforcement Learning (RL) framework as an alternative to supervised learning to train models. Similar to Vinyals et al. (Vinyals et al.,, 2015), they also employed the Pointer Network architecture and studied the TSP in their research. Kool et al. (Kool et al.,, 2018) applied a transformer architecture Vaswani et al., (2017), with a transformer encoder and a point network type of decoder to solve TSPs and vehicle routing problems (VRPs). Nazari et al. (Nazari et al.,, 2018) extended the VRP, modified the encoder by replacing the LSTM unit with 1D convolutional layers to have the model invariant to the order of input sequence, handling the dynamic state changes.
Subsequent studies followed the RL framework with different architectures and training methods to solve CO problems, mainly for TSPs and VRPS (Miki et al.,, 2018; Deudon et al.,, 2018; Ma et al.,, 2019; Kwon et al.,, 2020; Xu et al.,, 2021). Beginning with well-known TSP and VRP problms, research efforts have extended to the JSSP with a similar framework.
Zhang et al. Zhang et al., (2020) proposed a deep reinforcement learning (DRL) model to learn how to prioritize an operation within a possible operation set. They utilized a graph neural network to embed nodes (operations) in a disjunctive graph Adams et al., (1988) to represent states effectively. They designed the reward function as the quality difference between the partial solutions corresponding to the two states $s_{t}$ and $s_{t+1}$. Rather than employing the common makespan of a schedule as the quality metric, this approach utilizes an estimated completion time to measure the quality of a specific state $s_{t}$. Furthermore, the quality difference between the two states tends to promote a greedy selection by the temporal reward. However, opting for an action with a slightly lower immediate reward can often lead to a more favorable overall schedule. This limitation results in a relatively large optimality gap, $17.7\%$ in even small 6 machines $\times$ 6 jobs synthetic instances. To address these limitations, we use actual makespan as the reward metric and episodic reward instead of temporal reward to prevent greedy selection, We demonstrate better performance, $4.8\%$ optimality gap in $6\times 6$.
Numerous studies Park et al., 2021b ; Park et al., 2021a ; Chen et al., (2022); Yuan et al., (2023) have adopted a graph neural network or its variant to embed operations similar to the one proposed by Zhang et al. (Zhang et al.,, 2020) with different reward functions, training algorithms, masking technique and model structures. The graph neural network shares information between operations connected by conjunctive and disjunctive edges within a disjunctive graph, blending two types of information.
@@
Apart from using a graph neural network, Tassel et al. (Tassel et al.,, 2021) proposed an RL framework with seven designed scalers to capture the current state of individual jobs accurately. The major drawback of the proposed framework is problem size-dependent structure. They have to retrain or make adjustments to solve instances with different scales from the initial training. An advantage of our model is an agnostic problem size from an embedding scheme based on a modified transformer(Vaswani et al.,, 2017) structure.
Chien-Liang et al. Liu et al., (2020) introduced another RL framework based on a convolutional neural network (CNN) architecture. They utilize three matrices to represent states effectively. This study’s limitation lies in its relatively simple action space, primarily comprising several dispatching rules such as FIFO and SPT, which restrict candidates of the operations that can be scheduled and diminish solution quality.
@@ 3 Problem Definition
A Job Shop Scheduling Problem (JSSP) contains a set of $J$ jobs, for $j=1,2,\ldots,J$ which are to be processed in a shop that consists of $M$ machines for $m=1,2,\ldots,M$. Each job is comprised of $M$ operations, $O_{jm}$ for for $j=1,2,\ldots,J$ and $m=1,2,\ldots,M$ where each operation $O_{jm}$ can be processed on specific machine $m$. Additionally, each operation $O_{jm}$ has a predefined order $t$ in the sequence of operations. That is, $O^{t_{1}}_{jm}$ can be processed after all the prior operations, $O^{t}_{jm}$ where $t<t_{1}$, has been completed. One machine processes exactly one operation at a time without interruption. The time to complete a operation $O_{jm}$ is called the processing time, denoted as $p_{jm}$. Completion of all the element operations completes a job. A valid sequence of job assignments to machines generates a feasible schedule. The objective is to learn a schedule or a scheduling policy that optimizes a function of feasible schedules, e.g., the minimization of makespan or the maximization of machine utilization.
@@ 4 Attention-based Reinforcement Learning for JSSP
Our attention-based, reinforcement learning scheduler (ARLS) has an encoder-decoder structure that leverages feature embedding in a higher dimensional space. The input vector of each operation is encoded by linear projection, which is then transformed into embedding by the encoder using a self-attention mechanism. However, the encoder structure of ARLS is different from that of the original transformer (Vaswani et al.,, 2017).
@@ 4.1 Input
The input vectors $\mathbf{o}_{n}$, for $n=1,2,...,J\times M+1$, are concatenated from four vectors: 1) job index, 2) operation order in the job sequence, 3) machine index of the operation, and 4) processing time of the operation. The operation index $n=J\times M+1$ accounts for the cases when no operation is assigned (machine is idled).
The operations are initially encoded via the linear projection of the input as shown in 1.
@@ 4.2 Encoder
The encoder takes the encoding $\bm{h}^{(0)}$ as input and transforms it to the embedding $\bm{h}^{(L)}$ through multiple, e.g., $L$, encoder layers. We generally follow the structure of the encoder layers suggested by  (Vaswani et al.,, 2017) and introduce novel modifications for the JSSP. A masked, multi-head attention mechanism is utilized in the encoder, unlike the original transformer where simple multi-head attention is used in the encoder. The objective is to have attention better focused on associated operations, e.g., the operations in the same job sequence or the operations for the same machine.
@@
We apply two different masking schemes to $c_{i,j}$: one related to the jobs, denoted as $c_{i,j}^{\mathcal{(J)}}$, and another for the machines, $c_{i,j}^{\mathcal{(M)}}$, and integrate the two encoder results at the top of the stack. In particular, the values of the elements $c_{i,j}^{\mathcal{(J)}}=-\infty$ are masked if operations $i$ and $j$ are not in the same job sequence, and, similarly, $c_{i,j}^{\mathcal{(M)}}=-\infty$ if operations $i$ and $j$ are not processed by the same machine.
@@ 4.3 Decoder
The decoder has a single pillar of $L^{*}$ identical attention layers similar to the original transformer model  (Vaswani et al.,, 2017) However, masking is performed respective to the operations that were already assigned. Let $\mathcal{O}_{t}^{-}\subset\mathcal{O}$ be the set of operations that were already assigned by time $t$. The complement is $\mathcal{O}_{t}^{+}=\mathcal{O}\setminus\mathcal{O}_{t}^{-}$. Also, $\bm{h}_{t-1}$ denotes the encoder embedding of the operation assigned at the previous time step $t-1$.
@@ 4.4 Training
We modify the Monte-Carlo policy gradient method (Sutton and Barto,, 2018; Bahdanau et al.,, 2016) for training. The performance of learned policy was estimated by a Monte-Carlo method using observed makespans. The average of the makespans respective to the same problem is used as the baseline. Observed makespans minus the baseline, generate the advantages used for gradient updates
@@
where $\tau$ is a trajectory of an episode, $R(\tau)$ is the Monte-Carlo episodic reward (makespan of a trajectory), and $b$ is a baseline to reduce variance. The value from the value network is used as $b$ in many studies Mnih et al., (2013); Schulman et al., (2015); Haarnoja et al., (2018); Reda et al., (2020).
A different strategy compared to typical Monte-Carlo Policy Gradient generates multiple trajectories. A multi-trajectory strategy offers several advantages compared to employing a value network. More problem instances are expected to increase the scheduling performance robustness. And for each problem instance, the learner is expected to improve with more diverse scheduling solutions.
Let $\mathcal{T}$ denote the set of all trajectories, that is, all the feasible solutions for a JSSP instance. Let each $\tau_{n}\in\mathcal{T}$ be a sampled trajectory from time $t=0$ to $T$ where $T$ is the time a trajectory completes a schedule. For simplicity, we ignore the fact that $T$ can vary. and $\mathcal{T}=\langle\tau_{1},\ldots,\tau_{N}\rangle$ be the set of $N$ sampled trajectories for a JSSP instance. We sample $N$ parallel trajectories that enable us to produce a batch of disparate trajectories in parallel and enrich feasible solutions obtained for individual problem instances. One can control the sample size $N$. The equation 5 will be replaced with the following equation 6 for parallel trajectory.
@@ 5 Experimental Results
The training and testing approach for ARLS prioritizes rapid model learning in simplified settings. A key objective is a learner design which effectively scales to greater problem instance sizes and maintains high performance. ARLS is trained with synthetic data with 6 jobs and 6 machines. In total, 36 operations, with processing times randomly selected uniformly in the range between 1 and 15, similar to the prior study Zhang et al., (2020). The Adam optimization (Kingma and Ba,, 2014) was employed for weight updates. ARLS uses a relatively low complexity transformer architecture with only three encoder layers and a single decoder layer, the model dimension is 256, the number of heads is 16, and the fully connected feed-forward neural networks dimension is 512. Further design studies might improve performance.
Training generated 100,000 replicate synthetic 6 jobs $\times$ 6 machines problem instances. A mini-batch  (Yang et al.,, 2019) update is applied. The objective is to minimize the schedule’s makespan. To evaluate ARLS performance, the average optimality gap is used
@@
ARLS generalized performance is tested at varying sizes of synthetic and benchmark instances, including problem instance sizes greater than seen in training. To evaluate, we use three synthetic datasets with sizes $6\times 6$, $10\times 10$ and $15\times 15$. Additionally, seven well-known benchmark sets, from Shylo, (2010), are tested for performance.
Table 1 shows the test results on synthetic data. The average optimality gap is obtained from 1,000 instances for each problem size, which is compared with previous studies by Z20 (Zhang et al.,, 2020). Our model is trained on only $6\times 6$ setting. In contrast, Zhang et al., (Zhang et al.,, 2020) trained their model separately with samples from the three problem sizes. On average, ARLS outperforms their results by a substantial margin of 12.5%.
Table 1: AReL Performance on synthetic dataset
Jobs Machines Instances ARLS (Our) Z20
@@
ARLS was tested with seven benchmark datasets of varying sizes based on the number of jobs (ranging from 6 to 100) and machines (ranging from 5 to 20). Table 2 shows ARLS’s performance compared to work by Zhang et al., (2020) denoted as Z20, Park et al., 2021a as P21a, Park et al., 2021b as P21b, Chen et al., (2022) as C22, and Yuan et al., (2023) and Y23, and popularly adopted heuristic rules, such as first-in-first-out (FIFO), shortest processing time (SPT), and most work remaining (MWKR). Here, instead of the optimality gap, the best known gap, which is the relative difference to the best known makespan is used (as optimal makespans are unknown for some of the problems). Problems with multiple instances show average results. Our ARLS model showed the best performance on all but 4 problem sizes of 26 problem classes, despite being trained in the simplest setting of 6x6.
@@
Figure 1: Our ARLS model’s performance on benchmark datasets, ABZ, FT, YN, SWVM, and ORB
On average, ARLS outperformed other algorithms all cases, as shown in figures 1 and 2. Only in a few problem classes in TA was ARLS exceeded. Even then, the P21a algorithm was trained on problem instances more that 100% greater than the training size for ARLS.
Figure 2: Our ARLS model’s performance on benchmark datasets TA and DMU

### 18 Mar 2024 v2

@@ Abstract
Job shop scheduling problems represent a significant and complex facet of combinatorial optimization problems, which have traditionally been addressed through either exact or approximate solution methodologies. However, the practical application of these solutions is often challenged due to the complexity of real-world problems. Even when utilizing an approximate solution approach, the time required to identify a near-optimal solution can be prohibitively extensive, and the solutions derived are generally not applicable to new problems. This study proposes an innovative attention-based reinforcement learning method specifically designed for the category of job shop scheduling problems. This method integrates a policy gradient reinforcement learning approach with a modified transformer architecture. A key finding of this research is the ability of our trained learners within the proposed method to be repurposed for larger-scale problems that were not part of the initial training set. Furthermore, empirical evidence demonstrates that our approach surpasses the results of recent studies and outperforms commonly implemented heuristic rules. This suggests that our method offers a promising avenue for future research and practical application in the field of job shop scheduling problems.
@@ 1 Introduction
Job shop scheduling problem (JSSP) is a well-known and one of the hardest classes of combinatorial optimization problems in operations research, computer science, and industrial engineering. It has a long history as a research topic and is still being actively studied because of its applicability in numerous industrial areas, such as manufacturing and production domains (Gao et al.,, 2020).
@@
JSSP has been tackled in mainly two ways, by exact solution approaches and approximate approaches (Zhang et al.,, 2019; Kwon et al.,, 2021; Xie et al.,, 2019). Exact solution approaches with mathematical programming have been widely adopted (Ku and Beck,, 2016; Pei et al.,, 2020; Wang et al.,, 2018; Zhang and Wang,, 2018); however, it may take a prohibitive amount of time to find an optimal solution, or the problems can be infeasible to find an optimal solution. This can be problematic for real-world JSSP, where a large number of jobs or machines would be associated. This has led researchers and practitioners to turn their attention to approximate methods, which aim to achieve near-optimal solutions instead. Various heuristic approaches have been applied to solve JSSP (Gonçalves et al.,, 2005; Park et al.,, 2003; Peng et al.,, 2015; Kawaguchi and Fukuyama,, 2016; Chakraborty and Bhowmik,, 2015; Cruz-Chávez et al.,, 2017; Nagata and Ono,, 2018; Shi et al.,, 2020).
@@
Although the above mentioned approaches have gained success in addressing JSSP and are recognized as effective means to solve this particular problem type, solutions by those methods are subject to the constraints that are assumed when solving the problems. For example, solutions using the above methods are for the particular problems where the methods are applied, and thus, the solutions are not applicable to other problems in general. New problems are solved anew even with the same approach. Considering the shortcomings of classical solution approaches, it is essential to have a framework that can effectively solve large-scale problems without needing model adjustments or human intervention.
@@
To address those challenges, we propose a learning method that is capable of providing solutions to unseen problems. In particular, we employ attention mechanisms (Bahdanau et al.,, 2014) and integrate them with reinforcement learning (Sutton and Barto,, 2018) and the transformer architecture (Vaswani et al.,, 2017). We show that our trained learners can be reused to solve problems not only of the same size but also of larger sizes without retraining and demonstrate that our proposed approach outperforms the results in recent studies and widely adopted heuristic rules.
Building on this foundation, we highlight the principal contributions of our study:
  • The suggested model introduces a novel approach for representing the characteristics operations in JSSP by utilizing a revised transformer architecture, two independent transformer encoders attending operations in same job groups and machine sharing operations, respectively.
  • The proposed model is agnostic to problem size. A trained policy in simpler scenarios, like 6 jobs and 6 machines, can effectively solve unseen, more complex problem instances than used in training.
  • A trained policy with the proposed model outperforms popularly used dispatching heuristics and other proposed models employing RL frameworks.
  • We conducted extensive testing on synthetic and the seven most recognized benchmark instances of varying complexities to demonstrate the proposed model’s broad applicability and performance.
@@ 2 Related Work
In recent years, the application of machine learning and deep learning methods has expanded to the field of combinatorial optimization problems (COPs), providing innovative approaches to tackle problems that have traditionally been considered challenging due to their complexity and computational demands. Vinyals et al., (2015) proposed the Pointer Network, a sequence-to-sequence model to generate a permutation of the input sequence instead of a sequence from a fixed-size dictionary. This allows the model to select and order elements from the input sequence. In the original work of Vinyals et al., (2015), the traveling salesman problem (TSP) was studied. However, the model is trained with supervised learning and optimal solutions obtained from alternative methods such as the Held-Karp (Held and Karp,, 1962) algorithm for TSP. The study introduced the potential for implementing a learning-based approach in addressing COPs. A critical challenge, however, arises from its dependency on supervised learning, which requires optimal solutions for model training, but it is not possible to get optimal solutions for complex instances in most cases.
To address this issue, Bello et al., (2017) introduced a reinforcement learning (RL) framework as an alternative to supervised learning to train models. Like Vinyals et al., (2015), they also employed the Pointer Network architecture and studied TSPs in their research. The RL framework trains the model to learn an effective policy via a trial-and-error process, interacting with the problem environment to gradually improve its solution-finding capabilities. Unlike supervised learning used in Vinyals et al., (2015)’s approach, it does not require an optimal solution for model training, making it a practical approach for solving COPs.
Kool et al., (2018) applied a transformer architecture (Vaswani et al.,, 2017) with a transformer encoder and a point network type of decoder to solve TSPs and vehicle routing problems (VRPs). The proposed model, leveraging the Transformer architecture, captures the complex inter-relationships among inputs more effectively than the model by Bello et al., (2017), which employs Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber,, 1997). This advantage is attributed to the Transformer’s superior ability to represent long-range dependencies (Vaswani et al.,, 2017). Nazari et al., (2018) proposed a modified encoder architecture by replacing the LSTM unit with 1D convolutional layers to have a model that is invariant to the input sequence order, handling the dynamic state changes to solve extended the VRPs. Subsequent studies followed the RL framework with different architectures and training methods to solve combinatorial optimization COPs, mainly for TSPs and VRPs (Miki et al.,, 2018; Deudon et al.,, 2018; Ma et al.,, 2019; Kwon et al.,, 2020; Xu et al.,, 2021).
Beginning with well-known TSPs and VRPs, research efforts have extended to JSSP with a similar framework. Zhang et al., (2020) proposed a model that uses deep reinforcement learning to prioritize an operation from a set of possible operations. They used the Graph Neural Network (GNN) to represent operations’ states in a disjunctive graph. One advantage of using GNN to represent states is that the proposed model is agnostic to the size of the problem, i.e., the model will not be dependent on the number of jobs and machines of the problem instances. However, a drawback of their model lies in the reward function’s reliance on an estimated makespan instead of the actual makespan, potentially degrading the model’s performance.
Numerous studies Park et al., 2021a ; Park et al., 2021b ; Chen et al., (2022); Yuan et al., (2023) have adopted graph neural networks or variants to embed operations similarly to Zhang et al., (2020) with different reward functions, training algorithms, masking technique, and model structures. A motivation for using GNN is to represent the inter-relationships of operations in a disjunctive graph effectively. In a disjunctive graph, each operation is characterized by two main attributes: its relationship within a specific job group (via conjunctive edges) and its use of shared machines (via disjunctive edges). Potentially, keeping these attributes separate would help accurately represent the unique characteristics of each operation. However, a disjunctive graph is connected by both conjunctive and disjunctive edges, which results in a mixture of these important details.
@@
Apart from using a graph neural network, Tassel et al., (2021) proposed an RL framework with seven designed scalar variables to capture the current state of individual jobs. However, their framework is subject to the problem size and requires retraining or adjustments to the model to solve new problems of different sizes.
Liu et al., (2020) introduced another RL framework based on a convolutional neural network architecture for feature extraction. One notable drawback of the proposed model is employing too simple action space, selecting the next operation by three different dispatching rules such as first-in-first-out (FIFO), shortest processing time first (SPT), and longest processing time first (LPT). These simplistic action spaces could restrict candidates for the operations that can be scheduled and potentially degrade the quality of the solutions.
@@ 3 Problem Definition
A JSSP contains a set of $J$ jobs, for $j=1,2,\ldots,J$, which are to be processed in a shop that consists of $M$ machines for $m=1,2,\ldots,M$. Each job is comprised of $M$ operations, $O_{jm}$ for for $j=1,2,\ldots,J$ and $m=1,2,\ldots,M$, where each operation $O_{jm}$ can be processed on specific machine $m$. Additionally, each operation $O_{jm}$ has a predefined order $t$ in the sequence of operations. That is, $O^{t}_{jm}$ can be processed if and only if all the prior operations, $O^{t^{\prime}}_{jm}$, where $t^{\prime}<t$, have been completed. One machine processes exactly one operation at a time without interruption. The time to complete an operation, $O_{jm}$, is called the processing time, denoted as $p_{jm}$. Completion of all the element operations completes a job. A valid sequence of job assignments to machines generates a feasible schedule. The objective is to learn a schedule or a scheduling policy that optimizes a function of feasible schedules, e.g., the minimization of makespan or the maximization of machine utilization.
@@ 4 Attention-based Reinforcement Learning for JSSP
We propose a transformer-based reinforcement learning scheduler (ARLS) with enhancements on top of the original transformer architecture (Vaswani et al.,, 2017). ARLS specifically aims to capture the characteristics of operations better, focusing on job groups and machine-sharing aspects using two independent transformer encoders.
The input vector of each operation is encoded by linear projection, which is then transformed into embedding by the encoder using a self-attention mechanism. Two independent encoders produce two different operation embeddings from the same initial encoding. The decoder applies encoder-decoder attention using embeddings from both encoders and the decoder input. More details about this will be provided in subsequent sections.
@@ 4.1 Input
The input vectors $\mathbf{o}_{n}$, for $n=1,2,...,JM+1$, are concatenated from four vectors: 1) job index, 2) operation order in the job sequence, 3) machine index of the operation, and 4) processing time of the operation. The operation index $n=JM+1$ accounts for the cases when no operation is assigned. i.e., the machine is left idle.
The operations are initially encoded via the linear projection of the input as shown in equation 1.
@@ 4.2 Encoder
We have modified the original transformer encoder to effectively represent the characteristics of operations. In contrast to the traditional transformer encoder, ARLS has two separate encoders that both take the same initial input encoding, denoted as $\bm{h}^{(0)}$ as input. The input encodings are transformed from the initial encoding $\bm{h}^{(0)}$ to the embedding $\bm{h}^{(L)}$ through multiple, e.g., $L$, encoder layers.
The original transformer employs simple masked attention in its encoder layers to prevent the model from attending to padding, which is used to equalize the lengths of input tokens in natural language processing (NLP). In contrast to NLP, COPs have inputs of identical length, eliminating the need for padding. Instead, ARLS employs two distinct attention mechanisms in each encoder layer. These mechanisms are designed to focus attention specifically on operations within the same job group and operations sharing machines, respectively.
@@
We apply two different masking schemes to $c_{i,j}$, one for the jobs, denoted as $c_{i,j}^{\mathcal{(J)}}$, and another for the machines, $c_{i,j}^{\mathcal{(M)}}$, and integrate the two encoder results at the end of the encoder. In particular, the values of the elements $c_{i,j}^{\mathcal{(J)}}=-\infty$ are masked if operations $i$ and $j$ are not in the same job sequence, and similarly, $c_{i,j}^{\mathcal{(M)}}=-\infty$ if operations $i$ and $j$ are not processed by the same machine.
@@ 4.3 Decoder
The decoder has $L^{*}$ identical attention layers similar to the original transformer model (Vaswani et al.,, 2017). However, masking is performed respective to the operations that were already assigned. Let $\mathcal{O}_{t}^{-}\subset\mathcal{O}$ be the set of operations that were already assigned by time $t$. The complement is $\mathcal{O}_{t}^{+}=\mathcal{O}\setminus\mathcal{O}_{t}^{-}$. Also, $\bm{h}_{t-1}$ denotes the encoder embedding of the operation assigned at the previous time step $t-1$.
@@ 4.4 Training
We modify the Monte-Carlo policy gradient method (Sutton and Barto,, 2018; Bahdanau et al.,, 2016) for training. The performance of learned policy was estimated by a Monte-Carlo method using observed makespans. The average of the makespans respective to the same problem is used as the baseline. Observed makespans minus the baseline generate the advantages used for gradient updates
@@
where $\tau$ is a trajectory of an episode, $R(\tau)$ is the Monte-Carlo episodic reward, i.e., makespan of a trajectory, and $b$ is a baseline to reduce variance.
Preferred by researchers, as cited in works by (Mnih et al.,, 2013; Schulman et al.,, 2015; Haarnoja et al.,, 2018; Reda et al.,, 2020), the baseline $b$ is typically derived from the value network. Unlike the typical RL model, our proposed model does not have a value network but takes a different strategy compared to the typical Monte-Carlo Policy Gradient, which generates multiple trajectories. Our multi-trajectory strategy offers several advantages compared to employing a value network. More problem instances are expected to increase the scheduling performance robustness. Furthermore, for each problem instance, the learner is expected to improve with more diverse scheduling solutions.
Let $\mathcal{T}$ denote the set of all trajectories, that is, all the feasible solutions for a JSSP instance. Let each $\tau_{n}\in\mathcal{T}$ be a sampled trajectory, and $\mathcal{T}_{N}=\{\tau_{1},\ldots,\tau_{N}\}$ be the set of $N$ sampled trajectories for a JSSP instance, with which we make minibatch updates for the weights throughout the training as follows.
@@ 5 Experimental Results
The training and testing approach for ARLS prioritizes rapid model learning in simplified settings. A key objective is a learner design that effectively scales to greater problem instance sizes and maintains high performance. ARLS is trained with synthetic data with 6 jobs and 6 machines. In total 36 operations, with processing times randomly selected uniformly in the range between 1 and 15, similar to the prior study Zhang et al., (2020). The Adam optimization (Kingma and Ba,, 2014) was employed for weight updates. ARLS uses a relatively low complexity transformer architecture with only three encoder layers and a single decoder layer, the model dimension is 256, the number of heads is 16, and the fully connected feed-forward neural networks dimension is 512. Further design studies might improve performance.
Training generated 100,000 replicate synthetic 6 jobs $\times$ 6 machines problem instances. A mini-batch update (Yang et al.,, 2019) is applied. The objective is to minimize the schedule’s makespan. To evaluate ARLS performance, the average optimality gap is used
@@
The generalized performance of ARLS is tested with varying sizes of synthetic and benchmark instances, including problem instance sizes greater than seen in training. To evaluate, we use three synthetic datasets with sizes configured as follows: the first dataset has 6 jobs and 6 machines, the second dataset consists of 10 jobs and 10 machines, and the third dataset comprises 15 jobs and 15 machines. Additionally, seven well-known benchmark sets from Shylo, (2010) are tested for comparison.
Table 1 shows the test results on synthetic data. The average optimality gap is obtained from 1,000 instances for each problem size. To the best of our knowledge, the comparable test on synthetic instances is the study by Zhang et al., (2020)(Z20). Our model is trained on only 6 jobs and 6 machine settings. In contrast, Zhang et al., (Zhang et al.,, 2020) trained their model separately with samples from the three problem sizes. On average, ARLS outperforms their results by a substantial margin of 12.5%.
Table 1: Performance of our ARLS model compared with Z20 (Zhang et al.,, 2020) on three different synthetic datasets. Performance is evaluated by the optimality gap in percentage
Jobs Machines Instances ARLS (Ours) Z20
@@
ARLS was tested with seven benchmark datasets of varying sizes in the number of jobs (ranging from 6 to 100) and machines (ranging from 5 to 20). Table 2 shows ARLS’s performance compared to work by Zhang et al., (2020) denoted as Z20, Park et al., 2021a as P21a, Park et al., 2021b as P21b, Chen et al., (2022) as C22, and Yuan et al., (2023) and Y23, and popularly adopted heuristic rules, such as first-in-first-out (FIFO), shortest processing time (SPT), and most work remaining (MWKR). Here, instead of the optimality gap, the best-known gap, which is the relative difference to the best-known makespan, is used as optimal makespans are unknown for some of the problems. Problems with multiple instances show average results. Our ARLS model showed the best performance on all but 4 problem sizes out of 26 problem classes despite being trained in the simplest setting of 6 jobs and 6 machines.
@@
Figure 1: Generalized Performance of our ARLS model compared with heuristic rules and other studies on benchmark datasets, ABZ, FT, YN, SWVM, and ORB. Results by Zhang et al., (2020) denoted as Z20, Park et al., 2021a as P21a, Park et al., 2021b as P21b and Yuan et al., (2023) and Y23, and popularly adopted heuristic rules, such as first-in-first-out (FIFO), shortest processing time (SPT), and most work remaining (MWKR).
On average, ARLS outperformed other algorithms in all cases, as shown in Figures 1 and 2. ARLS was exceeded only in a few problem classes in TA. Even then, the P21a algorithm was trained on problem instances about 67% greater than the training size for ARLS.
Although ARLS was exceeded by P21a for three TA classes: TA(30, 15), TA(50, 20), and TA(100, 20) when ARLS trained in the 6 jobs and 6 machines scenario, ARLS outperformed P21a when trained on datasets of the equivalent size of P21a, resulting smaller optimality gaps, with 16.8% vs. 19.1% for TA(30,15), 9.96% vs. 13..5% for TA(50,20), and 4.4% vs. 6.7% for TA(100,20), respectively.
Figure 2: Generalized Performance of our ARLS model compared with heuristic rules and other studies on benchmark datasets, TA and DMU. Results by Zhang et al., (2020) denoted as Z20, Chen et al., (2022) as C22, and Yuan et al., (2023) and Y23, and popularly adopted heuristic rules, such as first-in-first-out (FIFO), shortest processing time (SPT), and most work remaining (MWKR).